For the following function, C computes a value, where if you add millions of dollars to the value, the result is the cost of implementing a city recycling project when x , as a percent (not its decimal equivalent), citizens participate.

FORMULA- C (x)= 1.5x / 100x
a) Using this model, approximately what is the cost if 60% of the citizens participate






Using this model, estimate the percentage of participation that can be expected if $5 million is spent on this recycling project? Set up an equation and solve algebraically. Round to the nearest whole percent.


Answer by stanbon(48535)

For the following function, C computes a value, where if you add millions of dollars to the value, the result is the cost of implementing a city recycling project when x , as a percent (not its decimal equivalent), citizens participate.
FORMULA- C (x)= 1.5x / 100x
a) Using this model, approximately what is the cost if 60% of the citizens participate
C(60) = 1.5*60/(100*60) = 90/6000 = 0.015 million dollars
==========================
Using this model, estimate the percentage of participation that can be expected if $5 million is spent on this recycling project? Set up an equation and solve algebraically. Round to the nearest whole percent.
5 = 1.5x/100x
500x = 1.5x
This is a contradiction unless x = 0.
I think your formula for C is wrong.
==============
Cheers,
Stan H.